A wire .50 m long carrying a current of 8 A is at an angle 15 degrees to a .40 T field. what is the magnetic field force excerted on the wire
1
Expert's answer
2012-05-31T09:18:19-0400
F = B * I * L * sinA. Given:
B = 0.40 T
I = 8 A
L = 0.50 m
A = 15
degrees
F = 0.40 * 8 * 0.50 * sin15 = 0.40 * 8 * 0.50 * 0.2588 = 0.4141 N.